# LLM Chain Ensembles for Scalable and Accurate Data Annotation

### v1

@@ II Related Works
Our proposed LLM ensemble method takes a similar design approach to FrugalGPT [24], which aligns LLMs in a chain and routes subsets of data that do not meet a predefined “reliability score” threshold to the next LLM in the chain. [^1] The authors refer to this sequential design as a LLM cascade instead of a chain. However, our solution differs significantly. First, their “reliability score” requires subsets of in-domain labeled data, whereas our LLM chain ensemble uses a generalized confidence score for zero-shot predictions, eliminating the need for any prior data labeling. Furthermore, while their method discards the labels and scores of prior LLM predictions, our approach retains all available scores in a rank-based ensemble to select the best label. Finally, our work uniquely focuses on scalability and data annotation for complex data domains and tasks presented by CSS, where traditional routing or ensemble methods may not be as effective.
@@ IV-B Misinformation Detection
Misinformation detection involves identifying ”false or inaccurate information that is deliberately created and intentionally or unintentionally propagated” [30]. In a corporate context, especially during a crisis, the ability to swiftly detect and mitigate misinformation is crucial for maintaining public trust and protecting brand reputation. Misinformation can lead to widespread public confusion, unfounded fear, or misguided actions that may have direct implications for a company’s products, services, or public image.
@@ IV-C Ideology
We define ideology as ”the shared framework of mental models that groups of individuals possess, providing both an interpretation of the environment and a prescription as to how that environment should be structured” [32]. In various contexts, including public health and other critical areas, understanding consumer ideology can be instrumental in tailoring communication to make quality information more accessible and persuasive. When messaging aligns with an individual’s ideological stance, it can enhance receptiveness and trust in the information source.

### v2

@@ II Related Works
Our proposed LLM ensemble method takes a similar design approach to FrugalGPT [24], which aligns LLMs in a chain and routes subsets of data that do not meet a predefined “reliability score“ threshold to the next LLM in the chain. [^1] The authors refer to this sequential design as a LLM cascade instead of a chain. However, our solution differs significantly. First, their “reliability score” requires subsets of in-domain labeled data, whereas our LLM chain ensemble uses a generalized confidence score for zero-shot predictions, eliminating the need for any prior data labeling. Furthermore, while their method discards the labels and scores of prior LLM predictions, our approach retains all available scores in a rank-based ensemble to select the best label. Finally, our work uniquely focuses on scalability and data annotation for complex data domains and tasks presented by CSS, where traditional routing or ensemble methods may not be as effective.
@@ IV-B Misinformation Detection
Misinformation detection involves identifying “false or inaccurate information that is deliberately created and intentionally or unintentionally propagated” [30]. In a corporate context, especially during a crisis, the ability to swiftly detect and mitigate misinformation is crucial for maintaining public trust and protecting brand reputation. Misinformation can lead to widespread public confusion, unfounded fear, or misguided actions that may have direct implications for a company’s products, services, or public image.
@@ IV-C Ideology
We define ideology as “the shared framework of mental models that groups of individuals possess, providing both an interpretation of the environment and a prescription as to how that environment should be structured” [32]. In various contexts, including public health and other critical areas, understanding consumer ideology can be instrumental in tailoring communication to make quality information more accessible and persuasive. When messaging aligns with an individual’s ideological stance, it can enhance receptiveness and trust in the information source.